OpenStudy (anonymous):

"Find x > 3 such that ln(x) < x^(0.1)" How do you solve this type of problem? I plugged it into my TI-89 solver and didn't get anything. As a hint, the number is unbelievably huge.

OpenStudy (anonymous):

my teacher gave us this problem in class today and no one knew what to do can someone help?

OpenStudy (anonymous):

seems hard can someone hekp me http://openstudy.com/study#/updates/4fc65a22e4b022f1e12dd2b9

OpenStudy (anonymous):

\[ \log x<x^{0.1}\\ \log_x\log x<0.1\\ \frac{\log\log x}{\log x}<0.1 \]There's a start, anyway.

OpenStudy (anonymous):

ok I have it but at least I know there is an answer now have to find out why x=8.8E+15 giving ln(x) = 36.71 x^(0.1) = 39.3

OpenStudy (anonymous):

@nbouscal the question was natural log not log so its to the base e not 10?

OpenStudy (anonymous):

When I say \(\log\) I mean the natural logarithm. If I mean \(\log_{10}\) I'd say it :P

OpenStudy (anonymous):

ok then say ln as that's what its called :P otherwise there is confusion!

OpenStudy (anonymous):

That's only what it's called in fields where using base 10 logarithms is useful in some way. In pure mathematics, 10 is just a number, it's not special. The default assumption should be the natural logarithm.

OpenStudy (anonymous):

Ah to change ln to log you divide by the conversion factor 2.303, might be good as you can't have ln with a base of x. the conversion to base 10 could be a start?

OpenStudy (anonymous):

Why would you convert to base 10? There's no mathematical reason to work in base 10. I already got rid of the logarithm with base x, when I simplified it to \[ \frac{\log\log x}{\log x}<0.1 \]

OpenStudy (anonymous):

ok, let me get what you did

OpenStudy (anonymous):

ok I see what you did now what

OpenStudy (anonymous):

Well, now we can subtitute \(t=\log x\), so we have \(\dfrac{\log t}{t}<0.1\)

OpenStudy (anonymous):

Solving \(\dfrac{\log t}{t}<0.1\), we learn that \(t>35.7715\), approximately. So we know that \(\log x >35.7715\). That means that \(x>e^{35.7715}\), or approximately \(x>3.43\times10^{15}\).

OpenStudy (anonymous):

That looks good, and its in the same ball park as my estimate. Nice job :D

OpenStudy (anonymous):

The only question is how we solve \(\log t<0.1t\). I used WolframAlpha, which in turn used the Lambert W-function, which is a relatively advanced technique. I'm sure we could have got there using a kind of guess and check, I'm just not sure whether that was the method expected by the teacher in this case.

OpenStudy (anonymous):

Sorry back now, so how did you get t>35.7715 ? was that using WolframAlpha?

OpenStudy (anonymous):

ok, got there using excel and seen it on Wolfram.hmmmm